OpenStudy (anonymous):

If f(x)=sum of (X^n)/(n!) for n=0 to infinity for all x, then f^(4)(0)=4! how do I figure this out?

myininaya (myininaya):

figure what out?

OpenStudy (anonymous):

if the answer is 4!

OpenStudy (anonymous):

if it's helps: lim (n->infinity)x^n/n! = 0 for every real number x

OpenStudy (anonymous):

not really but thanks, I'm reading through here, and it has to do with taylor series

OpenStudy (anonymous):

would it not be 1?

myininaya (myininaya):

f(x)=e^x f'(x)=e^x ... f''''(0)=e^0=1 thats what i think

OpenStudy (anonymous):

e^x at f''' x=0 yeah I do too thanks

OpenStudy (anonymous):

have another one

myininaya (myininaya):

what is it